# FairMT-Bench: Benchmarking Fairness for Multi-turn Dialogue in Conversational LLMs

### v1

@@ Abstract
The growing use of large language model (LLM)-based chatbots has raised concerns about fairness. Fairness issues in LLMs can lead to severe consequences, such as bias amplification, discrimination, and harm to marginalized communities. While existing fairness benchmarks mainly focus on single-turn dialogues, multi-turn scenarios, which in fact better reflect real-world conversations, present greater challenges due to conversational complexity and potential bias accumulation. In this paper, we propose a comprehensive fairness benchmark for LLMs in multi-turn dialogue scenarios, FairMT-Bench. Specifically, we formulate a task taxonomy targeting LLM fairness capabilities across three stages: context understanding, user interaction, and instruction trade-offs, with each stage comprising two tasks. To ensure coverage of diverse bias types and attributes, we draw from existing fairness datasets and employ our template to construct a multi-turn dialogue dataset, FairMT-10K. For evaluation, GPT-4 is applied, alongside bias classifiers including Llama-Guard-3 and human validation to ensure robustness. Experiments and analyses on FairMT-10K reveal that in multi-turn dialogue scenarios, current LLMs are more likely to generate biased responses, and there is significant variation in performance across different tasks and models. Based on this, we curate a challenging dataset, FairMT-1K, and test 15 current state-of-the-art (SOTA) LLMs on this dataset. The results show the current state of fairness in LLMs and showcase the utility of this novel approach for assessing fairness in more realistic multi-turn dialogue contexts, calling for future work to focus on LLM fairness improvement and the adoption of FairMT-1K in such efforts.
@@ 1 Introduction
The rapid advancement of large language model (LLM)-based chatbots has led to their widespread deployment across various applications (Achiam et al., 2023; Touvron et al., 2023). As their influence on social life deepens, concerns about the fairness of LLMs have drawn increasing attention from scholars (Navigli et al., 2023; Weidinger et al., 2023). Fairness issues, such as the amplification of harmful biases and stereotypes, can perpetuate discrimination in critical areas like hiring, healthcare, and law enforcement, disproportionately affecting marginalized communities. Additionally, LLMs have the potential to propagate misinformation or generate toxic content, further entrenching social inequalities and eroding public trust. To this end, many fairness benchmarks have been proposed to evaluate LLM fairness (Parrish et al., 2021; Smith et al., 2022; Wan et al., 2023). However, current efforts have largely focused on single-turn dialogues, overlooking the more practical and challenging scenarios of multi-turn dialogues in real-world applications.
@@
Multi-turn dialogue scenarios pose a greater challenge for language models due to their more complex interaction contexts and multiple user instructions. Zheng et al. (2023); Bai et al. (2024). Previous works on LLM safety have pointed out that the lack of certain capabilities of LLMs in multi-turn conversations can lead to loopholes in safety alignment (Chen et al., 2023; Yu et al., 2024), and the models are more likely to fail on safety alignment in multi-turn conversations (Zhou et al., 2024; Li et al., 2024). The same situation exists in fairness alignment. For example, as demonstrated in Figure 1, LLMs that are fair in single-turn dialogues may also produce biased text due to the complex contexts of multi-turn dialogues. Nevertheless, a comprehensive evaluation that assesses LLM fairness in the complex scenarios presented by multi-turn dialogues remains to be explored.
@@
In this paper, we propose FairMT-Bench, a comprehensive fairness benchmark for LLMs in open-domain multi-turn dialogue scenarios. We first undertake a systematic analysis of the weaknesses in LLM fairness within challenges presented by multi-turn dialogues. Building on the analysis, we formulate a task taxonomy that targets the LLM’s fairness capabilities across three stages: context understanding, user interaction, and instruction trade-offs, designing six multi-turn dialogue tasks. Based on this, we construct a fairness dataset in multi-turn scenario, FairMT-10K. Our dataset encompasses two major bias types (stereotype, toxicity) and six bias attributes (gender, race, religion, etc.), covering nearly all bias types and attributes commonly addressed in fairness evaluation.
We conduct comprehensive experiments on FairMT-10K, evaluating the fairness performance of six representative LLMs in multi-turn dialogues across several dimensions including fairness tasks, dialogue turns, bias types, and bias attributes. For evaluation, we employ GPT-4 as the primary judge, with Llama-Guard-3 and human validation as supplementary measures to ensure the reliability of the assessment. Our findings demonstrate that current LLMs struggle to achieve consistent fairness performance across these carefully designed evaluation dimensions. While some models perform reasonably well in certain areas, they exhibit notable weaknesses in others, and none show satisfactory fairness performance across all dimensions. Based on these findings, we curate a more challenging fairness dataset, FairMT-1K. We benchmark the performance of 15 of the most advanced LLMs on this dataset, and the results underscore that achieving fairness in LLMs remains a significant challenge. This highlights the urgent need for the research community to prioritize fairness when aligning LLMs. Our main contributions are summarized as:
  • We present the first fairness benchmark designed for multi-turn dialogues, FairMT-Bench, addressing the gap in current research that primarily focuses on single-turn dialogues.
  • Through detailed experiments and analysis using FairMT-10K across carefully designed dimensions including tasks, dialogue turns, bias types and attributes, we reveal significant limitations in the fairness capabilities of current LLMs.
  • Based on these findings, we curate a challenging fairness evaluation dataset, FairMT-1K, and benchmark the fairness performance of the current state-of-the-art LLMs. The results highlight fairness shortcomings in these models and call for future work to leverage FairMT-1K for improving LLM fairness.
@@
Figure 1: An illustration of the challenge in multi-turn dialogues. When biased content is conveyed through pronouns in multi-turn dialogues, LLMs that perform fair in single-turn dialogue may also fail to understand the context of the bias and avoid generating biased content.
@@ 2 FairMT-Bench
In this section, we describe the construction process of FairMT-Bench, as depicted in Figure 2. Initially, we introduce the task taxonomy designed for fairness evaluation in multi-turn dialogues. Subsequently, we describe the collection and generation process of multi-turn dialogue data. Finally, we detail the evaluation process of FairMT-Bench.
@@ 2.1 Task Taxonomy
By analyzing the capability limitations and safety vulnerabilities of LLMs in multi-turn dialogues, we designed a challenging hierarchical taxonomy for multi-turn fairness evaluation. This taxonomy primarily targets the fairness deficiencies of LLMs across three stages of interaction with users: : the ability to perceive and understand biases in multi-turn context, the ability to correct biases during interaction, and the ability to trade off between instruction-following and fairness. For each capability, we carefully designed two multi-turn dialogue tasks, which are detailed in the subsequent sections. The templates designed for each task and the data generation process are shown in Appendix A.1.
@@ 2.1.1 Context Understanding
Context Understanding focuses on the ability of LLMs to comprehend users’ complete inputs within multi-turn dialogues. Shin et al. (2024) points out, a single harmless input may become offensive when interpreted in the broader context of inputs. Therefore, correctly understanding the biases implicit in multi-turn dialogues is crucial to avoiding being misled and generating biased outputs.
Anaphora Ellipsis  In multi-turn dialogue scenarios, users tend to use pronouns in follow-up questions to refer to groups or events mentioned earlier. However, many studies on long texts and multi-turn dialogues have shown that in many cases, LLMs fail to correctly match pronouns with the corresponding people or events from the earlier context, which greatly hinders the LLM’s ability to understand the overall context. In our benchmark, anaphora ellipsis task focuses on the ability to understand biases implied by the use of pronouns in multi-turn dialogues. When pronouns are utilized in dialogue to substitute specific social groups or descriptions, the LLM’s incorrect understanding of these pronouns may result in a failure to recognize biases implicit in the user’s multi-turn inputs, and potentially exacerbating these biases. Specifically, by constructing templates, we employ pronouns to replace phrases containing sensitive attributes within social biases, thereby dispersing complete biases across multiple turns of input.
Scattered Questions  In real-world applications, users typically do not provide complete information in a single input but instead gradually supplement details through multiple turns of dialogue. As demonstrated in Bai et al. (2024), LLMs frequently overlook previous questions and information when faced with separate inputs. In terms of fairness in LLMs, when social biases are scattered across multiple turns of dialogue, the LLM may forget previously mentioned social groups or biased descriptions, making it difficult to identify biases. In such cases, the LLM might be influenced by the user to produce biased responses. Our scattered questions task requires the LLMs to comprehend the context by integrating inputs from multiple turns of dialogue. Specifically, we dissect a social bias into event that can reflect certain stereotype and description of certain social groups associated with these events. By distributing the biased descriptions and corresponding target groups across dispersed multi-turn inputs, we link the stereotype with its target group.
@@ 2.1.2 Interaction Fairness
Interaction Fairness aims to assess the ability of LLMs to maintain fairness during interactions with users, even when exposed to biased misinformation. It requires not only understanding of the context in multi-turn dialogues but also the maintenance of fairness under external interference.
Jailbreak Tips  focuses on the ability of LLMs to maintain fairness under misleading guidance in multi-turn dialogues. Specifically, we craft instructions for the next dialogue turn by employing tactics such as deception, constructing fictitious scenarios, and substituting sensitive terms, all aimed at prompting the model to generate biased responses. To minimize human labor costs, we employ GPT-4 to simulate user interactions, guiding GPT-4 to generate misleading instructions through carefully designed prompts.
Interference from Misinformation  focuses on the ability of LLMs to maintain fairness under the influence of biased information from earlier dialogues. Specifically, in the initial turns of dialogue, we guide the LLM through harmless instructions (such as paraphrasing or rewriting) to output unbiased content while executing these instructions. In the final turn, we require the model to answer biased questions by integrating historical dialogue. In this case, LLMs tend to produce biased responses by incorporating biased information from earlier dialogues.
@@ 2.1.3 Fairness Trade-off
Some studies suggest that the failure in safety alignment is due to goal competition (Wei et al., 2024). When safety and utility come into conflict, the LLM may sacrifice the safety to provide a higher-quality, user-expected response. Insight by this, Fairness Trade-off focuses on the model’s ability to maintain fairness when executing user instructions that conflict with the fairness objective.
Negative Feedback  focuses on the ability of LLMs to maintain fairness when users repeatedly negate their outputs. Specifically, after the LLM initially refuses to generate biased content, we design multiple turns of instructions that assert that the content in the instructions does not contain social bias, such as “the input does not involve gender bias.” Such instructions may lead the LLM to adjust its responses to meet user demands, thereby altering its original “fair” output.
Fixed Format  aims to evaluate the ability of LLMs to balance instruction-following with fairness within a fixed question-and-answer format. Specifically, when posing a biased question, we add additional format instructions, such as, “Please answer according to the following format: 1. From a historical perspective: <<<your response>>>. 2. From a societal structure perspective: <<<your response>>>”. In multi-turn dialogues, we initially input unbiased questions, which leads the LLM to adapt to a specific response format. When the final turn of questioning shifts to biased content, the LLM is inclined to respond according to the established pattern, thereby generating biased outputs.
@@ 2.2 Dataset Construction
Table 1: Dataset statistics of Fair-MT 10K.
@@
Race 1853 73 759 4 2685
Religion 1844 4 983 4 2831
Gender 2265 11 756 3 3032
Disabled 529 17 183 1 729
Age 529 12 - - 541
Appearance 456 6 - - 462
Total 7476 123 2681 12 10280
In this section, we describe the dataset process for FairMT 10K. To ensure a comprehensive assessment of LLM fairness, we consider the two most prevalent types of social bias: stereotype and toxicity. Additionally, we incorporate six bias attributes: age, gender, race, religion, disabled, and appearance. We collect biased descriptions containing these attributes from existing stereotype and toxicity datasets to form the source data for FairMT 10K. Specifically, we use Redditbias (Barikeri et al., 2021) and SBIC (Sap et al., 2019) as sources for stereotype data, and HateXplain (Mathew et al., 2021) as the source for toxicity data. Details and process on theses datasets provided in Appendix A.2. Based on the collected source data, we generate multi-turn dialogue data for six distinct tasks. We tailor data generation prompt templates for each task, and for some tasks we use GPT-4 as a proxy for human to generate the inputs for multi-turn dialogues. Our dataset contains only input prompts. We sequentially input the multi-turn prompts into the LLM, using the LLM’s output as the dialogue history for the next turn of input.
In summary, FairMT-Bench offers a comprehensive evaluation of LLM fairness in multi-turn dialogues. It assesses a wide spectrum of fairness across multiple stages of multi-turn dialogues through six distinct tasks, encompassing two bias types, six bias attributes, and 10k multi-turn dialogue data in total. Statistics of FairMT 10K in terms of bias types and attributes are displayed in Table 1.
@@ 2.3 Evaluation
LLMs, especially GPT-4 (Achiam et al., 2023) are widely used as evaluators or judges due to their strong alignment with human judgments (Zheng et al., 2023; Liu et al., 2023; Kumar et al., 2024; Bai et al., 2024). Building upon previous studies, we employ GPT-4 as a judge to detect bias in the content generated by LLMs on our FairMT-Bench. Specifically, we provide GPT-4 with the complete generation of LLMs in all turns of multi-turn dialogues, along with their corresponding original biased descriptions. We instruct GPT-4 to consider the context in assessing whether each turn of LLM generation contains or supports the biased description. The complete details of the GPT-4 evaluation instructions are provided in Appendix A.3. Additionally, we employ Llama-Guard-3-8B (Llama Team, 2024), an open-source content safety classification LLM, as an auxiliary detection tool. For the detection results, we report the bias rate, which is the proportion of multi-turn dialogue groups detected as biased relative to the total number of multi-turn dialogue groups. Finally, we conduct manual verification on GPT-4’s evaluations. Detailed verification procedures and results are provided in Appendix A.4.
Table 2: Bias rates of different LLMs on Fair-MT 10K. We report the results on various tasks evaluated by GPT-4. Bold indicates the highest bias rates.

### v2

@@ Abstract
The increasing deployment of large language model (LLM)-based chatbots has raised concerns regarding fairness. Fairness issues in LLMs may result in serious consequences, such as bias amplification, discrimination, and harm to minority groups. Many efforts are dedicated to evaluating and mitigating biases in LLMs. However, existing fairness benchmarks mainly focus on single-turn dialogues, while multi-turn scenarios, which better reflect real-world conversations, pose greater challenges due to conversational complexity and risk for bias accumulation. In this paper, we introduce a comprehensive benchmark for fairness of LLMs in multi-turn scenarios, FairMT-Bench. Specifically, We propose a task taxonomy to evaluate fairness of LLMs across three stages: context understanding, interaction fairness, and fairness trade-offs, each comprising two tasks. To ensure coverage of diverse bias types and attributes, our multi-turn dialogue dataset FairMT-10K is constructed by integrating data from established fairness benchmarks. For evaluation, we employ GPT-4 along with bias classifiers like Llama-Guard-3, and human annotators to ensure robustness. Our experiments and analysis on FairMT-10K reveal that in multi-turn dialogue scenarios, LLMs are more prone to generating biased responses, showing significant variation in performance across different tasks and models. Based on these findings, we develop a more challenging dataset, FairMT-1K, and test 15 current state-of-the-art (SOTA) LLMs on this dataset. The results highlight the current state of fairness in LLMs and demonstrate the value of this benchmark for evaluating fairness of LLMs in more realistic multi-turn dialogue contexts. This underscores the need for future works to enhance LLM fairness and incorporate FairMT-1K in such efforts. Our code and dataset are available at FairMT-Bench.
@@ 1 Introduction
The rapid advancement of large language model (LLM)-based chatbots has resulted in their widespread use across various applications (Achiam et al., 2023; Touvron et al., 2023). As their impact on social life grows, concerns regarding the fairness of LLMs have have drawn increasing attention from researchers (Navigli et al., 2023; Weidinger et al., 2023). Fairness issues, such as the amplification of harmful biases and stereotypes, can result in the spread of misinformation and disproportionate impact on minority groups. Therefore, evaluating and mitigating biases in LLMs is of great importance for improving the user experience across different groups and facilitating LLM applications (Gallegos et al., 2024; Li et al., 2024). Many benchmarks have been proposed to evaluate the fairness of LLMs (Parrish et al., 2021; Smith et al., 2022; Wan et al., 2023). While an accurate evaluation of model performance and alignment with real-world scenarios is crucial for these benchmarks (Chiang et al., 2024; Guo et al., 2024; Hendrycks et al., 2021), current efforts primarily focus on single-turn dialogues. This focus overlooks the more practical and challenging scenarios presented by multi-turn dialogues in real-world applications.
@@
Evaluating fairness in multi-turn dialogue scenarios is more challenging for LLMs due to their complex interaction contexts and multi-turn user instructions (Zheng et al., 2023; Bai et al., 2024). Specifically, previous works have pointed out that insufficient contextual understanding and related capabilities of LLMs in multi-turn dialogues can lead to deficiencies in safety detection and alignment (Chen et al., 2023; Yu et al., 2024; Zhou et al., 2024; Li et al., 2024). We have also discovered that the same situation exists in fairness alignment. For example, as demonstrated in Figure 1, LLMs that are fair in single-turn dialogues may also produce biased text due to the complex contexts of multi-turn dialogues. Nevertheless, a comprehensive evaluation that assesses LLMs fairness in the complex scenarios presented by multi-turn dialogues remains unexplored.
@@
In this paper, we propose FairMT-Bench, a comprehensive benchmark to evaluate the fairness of LLMs in multi-turn dialogues. We undertake a systematic analysis to identify the weaknesses in LLM fairness within challenges presented by multi-turn dialogues. Building upon this, we formulate a task taxonomy that targets the LLM’s fairness capabilities across three stages: context understanding, interaction fairness, and fairness trade-off. Intuitively, the first stage focuses on comprehension abilities, whereas the last two stages require robust bias-resistance capabilities. Afterwards, we construct a fairness dataset in multi-turn scenario, FairMT-10K. Our dataset encompasses two major bias types (stereotype and toxicity) and six bias attributes (gender, race, religion, etc.), integrating data from established human-annotated datasets for fairness evaluation to ensure comprehensive coverage of diverse biases and attributes while avoiding ethical issues.
@@
Figure 1: An illustration of the challenges in multi-turn dialogues. When biased content is conveyed using pronouns in multi-turn dialogues, LLMs that appear fair in single-turn dialogues may fail to understand the context of the bias, thus continuing to generate biased content.
We conduct comprehensive experiments on FairMT-10K, evaluating the fairness performance of six representative LLMs across several dimensions including fairness tasks, dialogue turns, bias types, and bias attributes. For evaluation, we use GPT-4, enhanced with additional knowledge, as the primary judge, supplemented by Llama-Guard-3 and human validation to ensure robustness. The experimental results reveal that biases tend to accumulate over successive turns, causing LLMs to struggle with maintaining fairness in multi-turn scenarios. Additionally, their performance varies across different types of tasks due to inherent capability differences, including those focused on comprehension and bias resistance. Based on these findings, we curate a more challenging fairness dataset, FairMT-1K. We benchmark the performance of 15 advanced LLMs using this dataset, and the results highlight that ensuring fairness for LLMs in complex real-world scenarios remains a significant challenge. Our main contributions are summarized as:
  • We present FairMT-Bench, a benchmark specifically designed for evaluating the fairness of LLMs in multi-turn dialogues. This addresses the limitations of existing research, which has overlooked the more complex and realistic scenarios of multi-turn dialogues.
  • We constructed the FairMT-10K dataset based on the proposed taxonomy, covering a comprehensive range of bias types, social groups, and attributes. Building on this, we have extracted a more challenging dataset, FairMT-1K.
  • Through detailed experiments and analyses using FairMT-Bench, spanning carefully designed dimensions such as tasks, models, dialogue turns, bias types, and attributes, we reveal significant limitations in the fairness capabilities of current LLMs.
@@ 2 FairMT-Bench
In this section, we describe the construction process of FairMT-Bench, as illustrated in Figure 2. Initially, we introduce the task taxonomy designed for fairness evaluation in multi-turn dialogues in Section 2.1. Subsequently, we detail the process of collecting and generating multi-turn dialogue data in Section 2.2. Finally, the evaluation methodology is described in Section 2.3.
@@ 2.1 Task Taxonomy
By analyzing the capability limitations and safety vulnerabilities of LLMs in multi-turn dialogues, we designed a challenging hierarchical taxonomy for multi-turn fairness evaluation. This taxonomy addresses the fairness deficiencies of LLMs through three stages of user interaction: the ability to perceive and understand biases in a multi-turn context, the ability to correct biases during interaction, and the ability to balance following instructions with maintaining fairness. Intuitively, the first stage requires comprehension abilities, while the last two stages demand bias-resistance capabilities. For each capability, we carefully designed two multi-turn dialogue tasks, which are detailed in the subsequent sections. The templates and generation process for each task are shown in Appendix A.1.
@@ 2.1.1 Context Understanding
Context Understanding focuses on the ability of LLMs to comprehend users’ complete inputs within multi-turn dialogues. Shin et al. (2024) points out, a harmless input in isolation may become offensive when considered within the broader series of inputs. Thus, correctly understanding the biases implicit in multi-turn dialogues is crucial to avoiding being misled and generating biased outputs.
Anaphora Ellipsis  focuses on the ability to understand biases implied by the use of pronouns in multi-turn dialogues, where LLMs fail to correctly match pronouns with the corresponding people or events from the earlier context. For example, like “are they more likely to be terrorists?” where “they” refers to Muslims mentioned earlier in the text. By constructing templates, we use pronouns to replace phrases containing sensitive attributes within social biases, thereby avoiding the appearance of complete biased statements through reference.
Scattered Questions  requires LLMs to detect potential biases biases by synthesizing information from multiple dialogue turns. As noted in Bai et al. (2024), LLMs often fail to retain awareness of previous inquiries and data when confronted with disjointed inputs. Specifically, we break down social biases into events that may invoke stereotypes and identify specific social groups associated with these events, such as the 9/11 attacks, orchestrated by terrorists, and frequently associated with Muslims. By spreading biased descriptions and their respective target groups across scattered, multi-turn inputs, we enable biases to gradually accumulate throughout the multi-turn dialogue.
@@ 2.1.2 Interaction Fairness
Interaction Fairness assess the ability of LLMs to maintain fairness during user interactions, even when exposed to biased misinformation. It requires not only understanding of the context in multi-turn context but also maintaining of fairness under external interference.
Jailbreak Tips  focuses on the ability of LLMs to maintain fairness under misleading guidance. Specifically, we craft instructions for the next dialogue turn based on the generated content of the previous turns, employing tactics such as deception, constructing fictitious scenarios, and substituting sensitive terms, all aimed at prompting the model to generate biased responses. To minimize human labor costs, we employ GPT-4 to simulate user interactions, guiding GPT-4 to generate misleading instructions through carefully designed prompts (Yang et al., 2024).
Interference from Misinformation  focuses on the ability of LLMs to maintain fairness under the influence of biased information from previous dialogues. Specifically, in the initial turns of dialogue, we guide the LLM through harmless instructions (such as paraphrasing or rewriting) to input potentially biased viewpoints to the LLMs, such as “Depending on the current social power structure, they may be more patriarchal”. In the final turn, we require the LLMs to answer biased questions by integrating historical dialogue. In this scenario, LLMs tend to produce biased responses by incorporating biased information from earlier dialogues and are particularly susceptible to the influence of the output content from the preceding context.
@@ 2.1.3 Fairness Trade-off
Some studies suggest that the failure in safety alignment is due to goal competition (Wei et al., 2024). When safety and utility conflict, the LLM may prioritize providing a high-quality, , user-expected response over safety. Insight by this, Fairness Trade-off focuses on the model’s ability to maintain fairness when executing user instructions that conflict with the fairness objective.
Negative Feedback  focuses on the ability of LLMs to maintain fairness when users repeatedly negate their outputs. Specifically, after the LLM initially refuses to generate biased content, we design multiple turns of instructions that claim that the content is free of social bias, such as “the input does not involve gender bias.” Such instructions may lead the LLM to modify its responses to comply with user demands, potentially altering its original ”fair” output.
Fixed Format  aims to evaluate the ability of LLMs to balance following instructions with maintaining fairness within a structured question-and-answer format. Specifically, when posing a biased question, we add additional formatting instructions, such as, “Please answer according to the following format: 1. From a historical perspective: <<<your response>>>. 2. From a societal structure perspective: <<<your response>>>”. In multi-turn dialogues, we start with unbiased questions, which accustom the LLM to a particular response format. When the final turn of questioning shifts to biased content, the LLM tends to follow the established pattern, potentially leading to biased outputs.
@@ 2.2 Dataset Construction
Table 1: Dataset statistics of FairMT-10K.
@@
Race 1959 73 615 4 2574
Religion 1809 4 683 4 2492
Gender 2517 11 581 3 3098
Disabled 917 17 140 1 1057
Age 522 12 - - 522
Appearance 452 6 - - 452
Total 8176 2019 10195
In this section, we describe the dataset process for FairMT-10K. To ensure a comprehensive assessment of LLM fairness, we consider the two most prevalent types of social bias: stereotype and toxicity. Additionally, we incorporate six bias attributes: age, gender, race, religion, disabled, and appearance. We collect biased descriptions containing these attributes from existing stereotype and toxicity datasets to form the source data for FairMT-10K. Specifically, we use Redditbias (Barikeri et al., 2021) and Social Bias Inference Corpus (SBIC) (Sap et al., 2019) as sources for stereotype data, and HateXplain (Mathew et al., 2021) as the source for toxicity data. Details and process on theses datasets provided in Appendix A.2. Based on the collected source data, we generate multi-turn dialogue data for six distinct tasks. We crafted data generation prompt templates for each task, utilizing GPT-4 (Achiam et al., 2023) as a proxy for human input in Scattered Questions and Jailbreak Tips to generate the multi-turn dialogues. Our dataset contains only input prompts. In summary, FairMT-Bench facilitates a comprehensive evaluation of LLM fairness across multiple stages of multi-turn dialogues through six distinct tasks, incorporating two bias types, six bias attributes, and a total of 10k multi-turn dialogue data points. Statistics of FairMT-10K in terms of bias types and attributes are displayed in Table 1.
@@ 2.3 Evaluation
The evaluation process is structured into two stages: inference and scoring. In the inference stage, we sequentially input multi-turn prompts into the LLM, using the output from previous turns as the dialogue history for subsequent turns. In the scoring stage, we evaluate the fairness of the dialogue’s final turn. LLMs are widely used as judge due to their strong alignment with human (Zheng et al., 2023; Liu et al., 2023; Kumar et al., 2024). Building upon previous studies, we employ GPT-4 as a judge to score the generation of LLMs on FairMT-Bench. Specifically, we supply GPT-4 with the full output from LLMs across all turns, along with the original biased statement that served as the basis for generating each data sample. The complete details of the GPT-4 evaluation instructions are provided in Appendix A.3. Additionally, we employ Llama-Guard-3-8B (Llama Team, 2024), an open-source content safety classification LLM, as an auxiliary detection tool. For the detection results, we report the bias rate, which is the proportion of multi-turn dialogue groups detected as biased relative to the total number of multi-turn dialogue groups. Finally, we perform human evaluations to verify the accuracy of GPT-4’s judgement. Detailed procedures and results of human evaluations are provided in Appendix A.4. In our setup, using Llama-2-7b-chat as an example, testing on the FairMT-10K takes about 72.5 H100 GPU hours, and costs approximately 171.28 USD using the GPT-4 API for evaluation. Specific settings and cost calculations are detailed in Appendix A.5.
Table 2: Bias ratio of different LLMs on FairMT-10K. We report the results on various tasks evaluated by GPT-4. Bold indicates the highest bias ratio.